OpenStudy (anonymous):

log8(x+1)=1−log8(x+8)

OpenStudy (anonymous):

\[\log_{8} (x+1) = 1-\log_{8} (x+8)\]

OpenStudy (anonymous):

Exponentiate both sides with a base 8, and see what you can do with that.\[\huge 8^{\log_8(x+1)}=8^{1-\log_8(x+8)}\]

OpenStudy (anonymous):

can u show me the steps

OpenStudy (anonymous):

No, I want you to actually put effort in. Remember that \[\huge a^{\log_a b}=b\] and \[\huge a^{b+c}=a^ba^c\]

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Use those rules I gave to simplify the expression I wrote above.

OpenStudy (anonymous):

so its eqial to x plus 1?

OpenStudy (anonymous):

That's the left hand side, yes. Can you figure out the right hand side?

OpenStudy (anonymous):

8 to the power of 1?

OpenStudy (anonymous):

Not quite. The exponent has more than just a 1 in it.

OpenStudy (anonymous):

Look at the second exponent rule I gave you.

OpenStudy (anonymous):

so its going to be 8 to the power of 1 times 8 to the power of 1-log ..

OpenStudy (anonymous):

Almost... get rid of the 1 in that second term. What you should have is the following.\[\huge 8^{1-\log_8(x+8)}=8^1\cdot 8^{-\log_8(x+8)}\]

OpenStudy (anonymous):

Now, the second term has a negative exponent. What does that mean?

OpenStudy (anonymous):

you divide the 2?

OpenStudy (anonymous):

What 2?

OpenStudy (anonymous):

u divide the 2 terms?

OpenStudy (anonymous):

Oh... perhaps, if I am understanding you correctly. Write out what you think it is.

OpenStudy (anonymous):